OpenStudy (anonymous):

Shelley invested some money in an account that will mature in one year. If Shelley withdraws the money after the maturity date, she will receive her initial investment plus 5%, but if she withdraws her money now, she will only be able to withdraw 94% of what she put into the account. Shelley put $316 into the account initially. How much of her initial investment will she lose if she withdraws her money now?

OpenStudy (kamille):

First of all we need to know how much she would get if she wait until maturity date 316*1.05=331,8$ she will get Then 331,8$ -100% x $ - 94 % so x= 331.8*94/100 x=311,89 at least you have to do this : 331.8-311.89= 19.91 So the answer is 19.91

OpenStudy (anonymous):

thank you ! :)

OpenStudy (kamille):

Welcome :]